# Dynamics of holographic steady flows near a first-order phase transition

### v1

@@ 3.2 Condensation nucleus
In order to reveal the nonlinear instability of metastable steady flow, we choose a steady state in the supercooled region as the initial data and then apply a disturbance parameterized by intensity. Similarly, there exists a threshold that triggers the phase transition. Figure 4 shows the real-time dynamics of the energy density of the system near the threshold of the disturbance intensity, from which it can be observed that, unlike the case of spinodal decomposition, a critical state appears in the dynamical intermediate process, and the final state of the evolution depends on the relative size of the actual disturbance intensity and the threshold. The entire dynamical process is similar to the phase transition of supercooled water induced by condensation nuclei. The initial disturbance needs to generate a local low-energy region that acts as a seed nucleus. When the disturbance intensity approaches the threshold, the seed nucleus will reshape and quickly converge to a critical nucleus configuration, which moves along the flow direction of the background due to non-zero momentum. The closer the disturbance intensity is to the threshold, the longer the critical nucleus exists in the dynamical process, exhibiting a critical behavior Chen:2022cwi . Finally, for the case where the disturbance intensity is slightly less than the threshold, the critical nucleus melts into the background and the system evolves back to a homogeneous steady flow. On the contrary, for the case where the disturbance intensity is slightly greater than the threshold, the critical nucleus gradually grows and expands to form a low-energy phase, indicating that the system successfully crosses the dynamical barrier and transitions to another local ground state, that is, the phase-separated state with non-zero momentum. Such results indicate that the metastable steady flow also suffers from the nonlinear instability, and there exists a moving critical nucleus that serves as a dynamical barrier to trigger the phase transition.
4 Inhomogeneous source
@@
from which, together with the second law of thermodynamics, we expect that the inhomogeneous scalar source can provide a damping mechanism for the moving system to decay from a steady state with $J\neq 0$ to a static state with $J=0$.
4.1 For steady flow
@@
In fact, the final state must be a static configuration with zero flow velocity. This conclusion is universal and does not depend on the specific form of the scalar source. The reason is that the time-independent inhomogeneous source makes the Killing vector (18) with non-zero velocity invalid, meaning $\vartheta[\phi_{0}]\neq 0$. For the final state of the evolution, the vector $\partial_{t}$ is the only candidate for the Killing vector, indicating a static configuration. That is to say, under the action of a time-independent inhomogeneous obstacle, whose local structure destroys the symmetry of the solution space, the moving system is dynamically unstable and will inevitably lose momentum and degenerate into a static system.
4.2 For moving domain
@@
In section 3, we obtain the phase-separated configuration with non-zero momentum, which appears as a high-energy moving domain flowing steadily at a constant velocity in a low-energy bath. If an inhomogeneous scalar source is imposed on such a steady-state geometry, there is no doubt that the moving domain will eventually cease and the system stabilizes to a static phase-separated configuration with a Killing vector $\partial_{t}$, as discussed previously. Nonetheless, the dynamical intermediate processes still deserve further investigation, especially the dynamical behavior of the moving domain under the action of an inhomogeneous scalar source. We find two characteristic dynamical patterns that the moving domains with different flow velocities exhibit when interacting with the obstacle.
@@
For the initial data, we use the phase-separated configuration with non-zero momentum obtained through spinodal decomposition as the background and then impose the inhomogeneous scalar source (20) at the center of the low-energy phase. Similar to the previous subsection, such an operation generates an energy well in the central region of the low-energy bath as a potential barrier. After a period of evolution, the moving high-energy domain will collide with it, allowing us to reveal the dynamical behavior of the moving domain when interacting with a fixed obstacle. As expected, numerical results show that such a dynamical process depends on the flow velocity of the moving domain and the amplitude $p$ of the external source, which characterizes the strength of the potential barrier. There are two dynamical patterns exhibited by the moving domain. For the case of low velocity, as the amplitude $p$ increases, the moving domain exhibits passing, pinning and rebounding behaviors in sequence. When flowing at a high velocity, there is an intermediate range of the amplitude $p$ that causes the moving domain to split. To this end, without loss of generality, we select a steady state each in the spinodal region on the steady flow branches with flow velocities $v=0.2$ and $0.4$. The selection rule for energy density requires that these steady flows with different velocities degenerate into the same static fluid with energy density $\epsilon=0.295$, located in the spinodal region of the static branch with $v=0$, after the corresponding Lorentz transformation. After perturbations, the velocities of the resulting moving domains are $v=0.18$ and $0.37$ respectively. For each case, we continuously adjust the amplitude $p$ to reveal the dynamical properties of the moving domain under interaction with obstacles of different strengths.
@@ 5 Conclusion
In addition, the steady flow also exhibits consistency in terms of thermodynamic stability and dynamical stability. On the one hand, a thermodynamically unstable steady state with a negative specific heat suffers from linear dynamical instability. Under an arbitrarily small perturbation, a dynamical transition will occur, leading to the emergence of a phase-separated configuration with non-zero momentum. For such a time-dependent inhomogeneous steady state, there exists a Killing vector $\partial_{t}-v\partial_{x}$ with a uniform flow velocity $v$. On the other hand, a thermodynamically metastable steady state is dynamical unstable at the nonlinear level. The occurrence of dynamical transition requires that the intensity of the disturbance exceeds a threshold to push the system across a critical state that acts as a dynamical barrier and then evolve to another local ground state, a moving phase-separated state. For the thermodynamically stable steady state, the system is dynamically stable and located in a global ground state.
@@
Finally and most importantly, by introducing an inhomogeneous scalar external source, the real-time dynamics of a moving holographic first-order phase transition system under the action of a fixed obstacle are revealed. Since the local structure of the scalar source invalidates the Killing vector $\partial_{t}-v\partial_{x}$ with non-zero velocity, the system keeps losing momentum during the dynamical process until it degenerates to a static state with a larger entropy and a Killing vector $\partial_{t}$. Such results show that in a closed system with fixed energy, there is a dynamical path that allows the system to evolve from a moving steady state to a static state, indicating the higher stability of the static state. In particular, due to the blocking effect of the obstacle on motion, the moving high-energy phase exhibits four characteristic dynamical behaviors when interacting with it, rebounding, pinning, passing through, and splitting, depending on the competitive relationship between the flow velocity of the moving domain and the strength of the obstacle. Such results provide an intuitive physical picture for experiments with strongly coupled fluids.
@@
The interaction between moving objects and obstacles is an interesting topic, especially for the domain walls which are crucial to the dynamical transport properties of the system. In our work, the domain wall of the effectively one-dimensional system is only a point, which makes the dynamical properties of the domain wall neglected during the interaction process. In the future, an important extension is to consider effectively two-dimensional or even three-dimensional transport systems, corresponding to linear and planar domain walls respectively. In that case, the domain wall will deform when interacting with an obstacle, further revealing the effects of surface tension and distortion energy on the dynamical properties of the transport system. In addition, it is worthwhile to further explore the dynamical properties of other strongly coupled transport systems, such as holographic superfluids and turbulence, under interaction with obstacles. Such studies will definitely provide even richer physical phenomena.
@@ Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under Grant Nos. 12035016, 12075026, 12275350, 12361141825 and 12375058 as well as the National Key Research and Development Program of China with Grant No. 2021YFC2203001.

### v3

@@ III.2 Condensation nucleus
In order to reveal the nonlinear instability of metastable steady flow, we choose a steady flow in the supercooled region as the initial data and then apply a disturbance parameterized by intensity. Similarly, there exists a threshold that triggers the phase transition. Figure 4 shows the real-time dynamics of the energy density of the system near the threshold of the disturbance intensity, from which it can be observed that, unlike the case of spinodal decomposition, a critical state appears in the dynamical intermediate process, and the final state of the evolution depends on the relative size of the actual disturbance intensity and the threshold. The entire dynamical process is similar to the phase transition of supercooled water induced by condensation nuclei. The initial disturbance needs to generate a local low-energy region that acts as a seed nucleus. When the disturbance intensity approaches the threshold, the seed nucleus will reshape and quickly converge to a critical nucleus configuration, which moves along the flow direction of the background due to non-zero momentum. The closer the disturbance intensity is to the threshold, the longer the critical nucleus exists in the dynamical process, exhibiting a critical behavior [22]. Finally, for the case where the disturbance intensity is slightly less than the threshold, the critical nucleus melts into the background and the system evolves back to a homogeneous steady flow. On the contrary, for the case where the disturbance intensity is slightly greater than the threshold, the critical nucleus gradually grows and expands to form a low-energy phase, indicating that the system successfully crosses the dynamical barrier and transitions to another local ground state, that is, the phase-separated state with non-zero momentum. Such results indicate that the metastable steady flow also suffers from the nonlinear instability, and there exists a moving critical nucleus that serves as a dynamical barrier to trigger the phase transition.
IV Inhomogeneous source
@@
from which, together with the second law of thermodynamics, we expect that the inhomogeneous scalar source can provide a damping mechanism for the moving system to decay from a steady flow or a inhomogeneous stationary state with $J\neq 0$ to a static state with $J=0$.
IV.1 For steady flow
@@
In fact, the final state must be a static configuration with zero flow velocity. This conclusion is universal and does not depend on the specific form of the scalar source. The reason is that the time-independent inhomogeneous source (20) makes the Killing vector (18) with non-zero velocity invalid, meaning $\vartheta[\phi_{0}]=-v\partial_{x}\phi_{0}\neq 0$. For the final state of the evolution, the vector $\vartheta$ in the zero velovity case, namely $\partial_{t}$, is the only candidate for the Killing vector, indicating a static configuration. That is to say, under the action of a time-independent inhomogeneous obstacle, whose local structure destroys the symmetry of the solution space, the moving system is dynamically unstable and will inevitably lose momentum and degenerate into a static system.
IV.2 For moving domain
@@
In section III, we obtain the phase-separated configuration with non-zero momentum, which appears as a high-energy moving domain flowing steadily at a constant velocity in a low-energy bath. If an inhomogeneous scalar source is imposed on such a steady-state geometry, there is no doubt that the moving domain will eventually cease and the system stabilizes to a static phase-separated configuration with a Killing vector $\partial_{t}$, as discussed previously. Nonetheless, the dynamical intermediate processes still deserve further investigation, especially the dynamical behavior of the moving domain under the action of an inhomogeneous scalar source. We find two characteristic dynamical patterns that the moving domains with different flow velocities exhibit when interacting with the obstacle.
@@
For the initial data, we use the phase-separated configuration with non-zero momentum obtained through spinodal decomposition as the background and then impose the inhomogeneous scalar source (20) at the center of the low-energy phase. Similar to the previous subsection, such an operation generates an energy well in the central region of the low-energy bath as a potential barrier. After a period of evolution, the moving high-energy domain will collide with it, allowing us to reveal the dynamical behavior of the moving domain when interacting with a fixed obstacle. As expected, numerical results show that such a dynamical process depends on the flow velocity of the moving domain and the amplitude $p$ of the external source, which characterizes the strength of the potential barrier. There are two dynamical patterns exhibited by the moving domain. For the case of low velocity, as the amplitude $p$ increases, the moving domain exhibits passing, pinning and rebounding behaviors in sequence. When flowing at a high velocity, there is an intermediate range of the amplitude $p$ that causes the moving domain to split. To this end, without loss of generality, we select a steady flow each in the spinodal region on the branches with flow velocities $v=0.2$ and $0.4$. The selection rule for energy density requires that these steady flows with different velocities degenerate into the same static fluid with energy density $\epsilon=0.295$, located in the spinodal region of the static branch with $v=0$, after the corresponding Lorentz transformation. After perturbations, the velocities of the resulting moving domains are $v=0.18$ and $0.37$ respectively. For each case, we continuously adjust the amplitude $p$ to reveal the dynamical properties of the moving domain under interaction with obstacles of different strengths.
@@ V Conclusion
In addition, the steady flow also exhibits consistency in terms of thermodynamic stability and dynamical stability. On the one hand, a thermodynamically unstable steady flow with a negative specific heat suffers from linear dynamical instability. Under an arbitrarily small perturbation, a dynamical transition will occur, leading to the emergence of a phase-separated configuration with non-zero momentum. For such a time-dependent inhomogeneous stationary state, there exists a Killing vector $\partial_{t}-v\partial_{x}$ with a uniform flow velocity $v$. On the other hand, a thermodynamically metastable steady flow is dynamical unstable at the nonlinear level. The occurrence of dynamical transition requires that the intensity of the disturbance exceeds a threshold to push the system across a critical state that acts as a dynamical barrier and then evolve to another local ground state, a moving phase-separated state. For the thermodynamically stable steady flow, the system is dynamically stable and located in a global ground state.
@@
Finally and most importantly, by introducing an inhomogeneous scalar external source, the real-time dynamics of a moving holographic first-order phase transition system under the action of a fixed obstacle are revealed. Since the local structure of the scalar source invalidates the Killing vector $\partial_{t}-v\partial_{x}$ with non-zero velocity, the system keeps losing momentum during the dynamical process until it degenerates to a static state with a larger entropy and a Killing vector $\partial_{t}$. Such results show that in a closed system with fixed energy, there is a dynamical path that allows the system to evolve from a moving stationary state to a static state, indicating the higher stability of the static state. In particular, due to the blocking effect of the obstacle on motion, the moving high-energy phase exhibits four characteristic dynamical behaviors when interacting with it, rebounding, pinning, passing through, and splitting, depending on the competitive relationship between the flow velocity of the moving domain and the strength of the obstacle. Such results provide an intuitive physical picture for experiments with strongly coupled fluids.
@@
The interaction between moving objects and obstacles is an interesting topic, especially for the domain walls which are crucial to the dynamical transport properties of the system. In our work, the domain wall of the effectively one-dimensional system is only a point, which makes the dynamical properties of the domain wall neglected during the interaction process. In the future, an important extension is to consider effectively two-dimensional or even three-dimensional transport systems, corresponding to linear and planar domain walls respectively. In that case, the domain wall will deform when interacting with an obstacle, further revealing the effects of surface tension and distortion energy on the dynamical properties of the transport system. In addition, it is worthwhile to further explore the dynamical properties of other strongly coupled transport systems, such as holographic superfluids and turbulence, under interaction with obstacles. On the other hand, the boundary effects ignored in our work are also an interesting topic. Through different boundary conditions, we can simulate the interaction between the physical system and the external environment. At this time, the physical quantities of the system may not be conserved, such as energy and momentum, which is different from the closed system in this work. For an open boundary, the physical phenomena may change significantly, for example, the energy and momentum of a steady flow within a fixed interval may be continuously lost. Such a result will lead to a continuous weakening of the interaction between the steady flow and the obstacle, thereby changing the dynamic behavior of the system. Such studies will definitely provide even richer physical phenomena.
@@ Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under Grant Nos. 12035016, 12075026, 12275350, 12361141825, 12375058 and 12447129 as well as the National Key Research and Development Program of China with Grant No. 2021YFC2203001, and the China Postdoctoral Science Foundation with Grant No. 2024M760691.